OpenStudy (roadjester):

19. Electrons are ejected from a metallic surface with speeds of up to 4.60 X 10^5 m/s when light with a wavelength of 625 nm is used. (a) What is the work function of the surface? (b) What is the cutoff frequency for this surface?

OpenStudy (roadjester):

@wolfe8

OpenStudy (wolfe8):

I believe we use E=K-W K=1/2mv^2 and E=hf

OpenStudy (wolfe8):

f=v/wavelength

OpenStudy (roadjester):

For part a; i managed to get that. \[\large K_{max}= hf-\phi\]

OpenStudy (roadjester):

Where phi is the work function. The cutoff frequency is defined as \[\large \lambda_c= {hf \over \phi}\] But I can't get the right numerical value.

OpenStudy (wolfe8):

Yea I was supposed to put + instead of - there. That looks right. What do you mean right numerical value?

OpenStudy (roadjester):

oops \[hc\over \phi\]

OpenStudy (roadjester):

oh well, according to my book, phi=1.38 eV which i got. But i can't seem to get the cutoff frequency of lambda_c=3.34 X10^14Hz

OpenStudy (wolfe8):

Hmmmmmm

OpenStudy (wolfe8):

Ah we used the wrong Planck's constant. Because our energy for phi is in eV, we have to use 4.13×10^(−15) eV/s for h.

OpenStudy (wolfe8):

I meant eVs

OpenStudy (roadjester):

hmm, nope phi = 1.38 eV so the calculation would be 4.13 X10^-15 *.X10^8 / 1.38

OpenStudy (roadjester):

that doesn't give me 3.34X10^14

OpenStudy (anonymous):

Ok. just use Einstiens equation since speed is given, u know what is the k.e energy of the electron and input energy can be found out from wavelength Einstien says Input energy = work function + K.E tada.. :D

OpenStudy (anonymous):

is it 1.39 eV ?

OpenStudy (roadjester):

@mashy phi is 1.38 eV; I'm looking for the frequency which is 3.34 X10^14 Hz I CAN'T get 3.34 X10^14 Hz and I don't know how to get it.

OpenStudy (anonymous):

work function = h (threshold frequency) [S.I. units though]

OpenStudy (anonymous):

i get exactly that value 3.34 * 10^14

OpenStudy (anonymous):

i think u forgot to convert from eV to Joules.. work function = 1.38 * 1.6* 10^-19

OpenStudy (anonymous):

Cut off frequency should be as follows:\[\phi=hf _{c}\]\[f _{c}=\phi/h\]Using that and the Planck's constant in EV-s (4.136e-15), I get a cutoff frequency of 3.34e14 (898.4nm).

OpenStudy (roadjester):

How did @PsiSquared derive the equation for the cutoff frequency? @douglaswinslowcooper @wolfe8 @ybarrap

OpenStudy (anonymous):

There is a minimum amount of energy needed to escape, Emin = h f, and phi is the same as Emin.

OpenStudy (anonymous):

derived the cutoff frequency just as @douglaswinslowcooper suggested.

OpenStudy (roadjester):

could you elaborate @PsiSquared

OpenStudy (anonymous):

Given:\[hf=\phi+E _{K}\]you want to find the frequency of light required to just liberate an electron without giving it any additional energy. The only reason that electron's liberated by the photoelectric effect have kinetic energy is because they're given energy beyond what is needed to free them. If the kinetic energy is zero, then the above equation becomes\[hf=\phi \]This equation describes the energy needed to counter the energy binding the electron to the surface. An electron with that energy would be free but have no kinetic energy. The work function, φ, tells just tells us the minimum energy required to free said electron. Does that make sense?

OpenStudy (roadjester):

yup; thanks!